Find the perimeter of a square with the dimensions 4x. What is the length of the square if its area is 256 sq.cm?
Answer by checkley77(12844)

Area=lw
Area=x*x=x^2 ans.
256=x^2
x^2=256
x=sqrt256
x=16 cm. is the length of the sides.